OpenStudy (anonymous):

determine whether 3x-5y=11 and 15x+9y=10 are parallel perpendicular or neither

OpenStudy (anonymous):

if slopes are equal ,lines are parallel. if their product is -1,lines are perpendicular.

OpenStudy (victoriasushchik):

14y + 5x + 7 = 5 + 5x Subtract 5x on both sides 14y+7=5 Subtract 5 on both sides 14y=-2 Divide 14 on both sides y=-1/7 So the slope is 0 and the y intercept is -1/7 2)3x - 5y + 5 = 0 Subtract 3x on both sides -5y+5=-3x Subtract 5 on both sides -5y=-3x-5 Divide -5 to everything y=(3/5)x+1 So slope is 3/5 and y-intercept is 1 3) 3x - 4y = -15 8x + 6y = -15 Put them both in slope-intercept form. Parallel lines have same slope, perpendicular lines have negative reciprocal slopes, and neither has neither. 3x-4y=-15 -4y=-3x-15 y=(3/4)x+(15/4) Slope is 3/4 8x+6y=-15 6y=-8x-15 y=(-4/3)x-(5/2) Slope is -4/3. So they are perpendicular 4)3x - 4y = -8 8x + 6y = -8 3x-4y=-8 -4y=-3x-8 y=(3/4)x+2 8x+6y=-8 6y=-8x-8 y=(-4/3)x-(4/3) They are perpendicular 5) y + 18 = -4x 6y = 36x - 12 y+18=-4x y=-4x-18 6y=36x-12 y=6x-2 Neither 6)4x - 2y = -15 4x + 3y = -15 4x-2y=-15 -2y=-4x-15 y=2x+(15/2) 4x+3y=-15 3y=-4x-15 y=(-4/3)x-5 neither 7)12x + 4y = 16 15x + 5y = 24 12x+4y=16 4y=-12x+16 y=-3x+4 15x+5y=24 5y=-15x+24 y=-3x+(24/5) Parallel 8) 9x + 3y = 12 12x + 4y = 19 3y=-9x+12 y=-3x+4 4y=-12x+19 y=-3x+(19/4) Parallel 9) 3x=-8 9y=-8 x=-8/3 Slope is undefined y=-8/9 Slope is 0 Perpendicular 10) (-4, -8) and (4, 1) Find slope. Slope is change in y divided by change in x (1--8)/(4--4) 9/8 Put in slope-intercept form y=mx+b y=(9/8)x+b To find b, substitute in one of the ordered pairs -8=(9/8)(-4)+b -8=-9/2+b -7/2=b So: y=(9/8)x-(7/2)

OpenStudy (anonymous):

means product of slopes